OpenStudy (anonymous):

What is the solution to the equation? 4 + 3s – 9s = 22 A. s = –3 B. s = –1.5 C. s = 1.5 D. s = 4 2/3

OpenStudy (anonymous):

A

OpenStudy (arabpride):

4 + 3s - 9s = 22 combine like terms: 3s - 9s + 4 = 22 -6s + 4 = 22 -6s + 4 = 22 -4 -4 -6s = 18 /-6 /-6 s = 3 Your correct answer is A: s = 3 *^.^*

OpenStudy (anonymous):

yay